OpenStudy (anonymous):

Expand log 2 (8/5(4)) 6 Log base 2 parentheses 8 over 5 power 4 parentheses to the power of 6 ?

OpenStudy (anonymous):

\[\large \log_2\left(\frac{8}{5^4}\right)^6\]

OpenStudy (anonymous):

start with \[6\log_2(\frac{8}{5^4})\]

OpenStudy (anonymous):

then turn the division in to a subtraction \[6\left(\log_2(8)-\log_2(5^4)\right)\]

OpenStudy (anonymous):

bring out the exponent of 4 \[6\left(\log_2(8)-4\log_2(4)\right)\]

OpenStudy (anonymous):

and finally see if you know what \(\log_2(8)\) is

OpenStudy (anonymous):

Ohhh okay which is 3